Explain why the quotient of x^8 and x^2 is x^6 instead of x^4.
Answer by nyc_function(2733)

When dividing one term with an exponent by another term having an exponent, exponents are subtracted.
So, x^8/x^2 = x^(8-2) = x^6
Understand?